quadratic equation within word problem.
A fenced in rectangular region with the dimensions 6 yards by 8 yards is suddenly expanded by the same distance in each dimension. The resulting region is now 143 square yards in size. How much distance was added to each dimension?
_____yards

Answer by stanbon(75887)

A fenced in rectangular region with the dimensions 6 yards by 8 yards is suddenly expanded by the same distance in each dimension. The resulting region is now 143 square yards in size. How much distance was added to each dimension?
_____yards
----------------------
Equation:
(6+x)(8+x) = 143 sq yds
----
x^2 + 14x + 48 = 143
---
x^2 + 14x - 95 = 0
----
Positive solution:
Ans: x = 5 yds (amt each dimension must be increased)
------------------
Cheers,
Stan H.
---------------